Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

You are wearing a pair of cargo pants with six pockets. You've put $10 in one of the pockets, but you cannot remember which one. After checking two pockets without success, what is the probability that the money will be in the next pocket you check?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

it's 1/4 right?

OpenStudy (anonymous):

i got it haha

OpenStudy (anonymous):

There are 6 pockets all together. You checked two of them and it wasn't there. So you can forget about those 2, and now there are 4 pockets left. The total number of ways for your next try is 4. The number that will be successful is 1. The probability of success on the next one is 1/4 = 25% . ======================================================= Another method of solution, somewhat more complex: Total number of pockets = 6. Amount of money missing = $10 First partial sum, i.e. sum of numbers in the first part of the problem = 6 + 10 = 16. Total number of pockets = 6 Current value of cargo pants on a fashion scale from 1 to 10 = -2. Second partial sum = 6 - 2 = 4. Probability = (second partial sum) / (first partial sum) = 4/16 = 25% .

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
10 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
13 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!